OpenStudy (anonymous):

Solve the following equation: x(13 − x) = 42

OpenStudy (anonymous):

first, x(13) is what?

OpenStudy (anonymous):

13x?

OpenStudy (anonymous):

x13-x^2=42

OpenStudy (anonymous):

correct, and x(-x) is what?

OpenStudy (anonymous):

-x^2

OpenStudy (anonymous):

-x^2+13x-42=0 x^2-13x+42=0

OpenStudy (anonymous):

so, I put it together like Irodd said, then...

OpenStudy (anonymous):

(x-7)(x-6)=0

OpenStudy (anonymous):

i put it in order...

OpenStudy (anonymous):

yes! now we have 13x-x^2=42

OpenStudy (anonymous):

x=7 and x=6

OpenStudy (anonymous):

x=7, x=6

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

he did it and gave you the answer... so I tried...

OpenStudy (anonymous):

Thanks everyone

OpenStudy (anonymous):

no prob.

OpenStudy (anonymous):

you still helped, it's ok lol next question?

OpenStudy (anonymous):

sure.